第 7 回ゼータ若手研究集会 講演アブストラクト

2 月 14 日 (金)

Ade Irma Suriajaya On the Zeros of the k-th Derivative of the Riemann Zeta Function under the Riemann Hypothesis

The number of zeros and the distribution of the real part of non-real zeros of the derivatives of the Riemann zeta function have been investigated by Berndt, Levinson, Montgomery, and Akatsuka. Berndt, Levinson, and Montgomery investigated the general case, meanwhile Akatsuka gave sharper estimates for the first derivative of the Riemann zeta function under the truth of the Riemann hypothesis. In this talk, we shall introduce the generalization of the results of Akatsuka to the k-th derivative (for positive integer k) of the Riemann zeta function.

李正勲 On the Artin-Mazur Zeta Functions of Rational Maps over ℂ_{p}

The Artin-Mazur zeta functions, which are sometimes referred as the dynamical zeta functions, were introduced by M. Artin and B. Mazur in 1965. The theory on these zeta functions in dynamical systems has then been developed. In this talk, we will focus on the result on the Artin-Mazur zeta functions of rational maps over ℂ_{p}. We will try to understand the statement and the proof of it. At the end of this talk, we will see an open problem suggested by A. Bridy. This research is motivated by A. Hinkkanen’s result on the Artin-Mazur zeta functions of rational maps over ℂ.

平島拓真 Newton polygon を用いた二次体の類数の可除性問題について

2 次体の類数の可除性問題とは，整数 n で類数の割れるものが無限に存在するかというもので，任意の n について肯定的に解かれている．さらに， Gauss の種の理論により， n = 2 の場合は，類数の可除性について必要十分条件が古典的に知られている． Kishi-Miyake (2000) は，類数の n = 3 で割れる 2 次体の必要十分条件を述べており， Newton polygon を用いて考察がなされている．本講演では，この Newton polygon を用いて得られた結果について報告する．平井卓哉多重フィボナッチゼータ関数

リーマンゼータ関数の n の部分をフィボナッチ数列 f_{n} に置き換えたものをフィボナッチゼータ関数という．フィボナッチゼータ関数をうまく多重化することで，リーマンゼータ関数と関連性があることが分かった．今回は，この関連性と多重フィボナッチゼータ関数のいくつかの性質について紹介する．

石川勲 p 進上半平面上の乗法的積分と肥田 p 進族

重さ 2 の Hilbert 固有形式に付随する肥田 p 進族から定まる 2 変数 p 進 L 関数が構成されるが，特に保型形式が楕円曲線からくる場合，その 2 変数 L 関数の critical line 上微分値が Heegner point のある p 進的な log での値であることが Darmon および Mok らによって示されていた．本講演ではこのことをアーベル多様体に一般化する過程において，ある p 進 L 関数の親戚のようなものと幾何的な点を結びつける公式を得たのでそれらについて話をしたい．

2 月 15 日 (土)

Kyuwool Han On the class number for imaginary quadratic number fields

We will consider about the class number one problem for imaginary quadratic number fields.

杉山真吾相対跡公式と Hilbert モジュラー形式の L 関数の劣凸評価

Ramakrishnan と Rogawski は，素数レベルの正則楕円尖点形式の虚 2 次ベースチェンジに付随する L 関数の中心値の平均の，レベルを無限大にするときの漸近公式を得た． Feigon と Whitehouse は彼らの結果を，代数体が総虚 2 次拡大の時における square free レベルの Hilbert 尖点形式の場合に拡張した．この講演では彼らの漸近公式の，総虚 2 次拡大という条件を外した上での一般のレベルの場合への拡張について話す．またこの公式の応用として， Hilbert 尖点形式に対する 2 次ベースチェンジ L 関数の重さアスペクトに関する劣凸評価を紹介する．これは都築正男氏 (上智大学) との共同研究である．

Relative trace formulas and subconvexity estimates of L-functions for Hilbert modular forms

Ramakrishnan and Rogawski gave an asymptotic formula for a mean of central L-values attached to the imaginary quadratic base change of elliptic holomorphic cusp forms with prime level as the level tends to infinity. Feigon and Whitehouse extended their result to the case of Hilbert cusp forms with square free level when the quadratic extention involved is totally imaginary. In this talk, we report our recent generalization of their asymptotic formula for Hilbert cusp forms to arbitrary levels, dropping the totally imaginary condition of the quadratic extension at the same time. As an application of this, we obtain a subconvexity bound of quadratic base change L-functions for Hilbert cusp forms in the weight aspect. This is a joint work with Masao Tsuzuki (Sophia University).

山本修司 2-labeled poset に付随する積分と多重ゼータ値

Euler-Zagier 型の多重ゼータ値に対する反復積分表示はよく知られており，その応用も豊富である．この講演では，有限半順序集合で記述される積分領域を考えることで上述の積分を一般化する．また応用として，等号付き多重ゼータ値のいくつかの関係式が簡潔に証明できる例を紹介する．塩見大輔 Ordinary でない円分関数体の無限族の構成

昨年の若手ゼータでは q(=係数体の位数) が素数でない場合に，円分関数体が ordinary となるための同値条件を与えた．本講演では， q が素数の場合に ordinary 性の議論をする．特に， Carlitz-Bernoulli 多項式を用いることで， ordinary でないような円分関数体 (導種は既約) の無限族が構成できることを紹介する．

跡部発 Pullbacks of hermitian Maass lifts

Hermitian Maass lift とは，楕円保型形式から (次数 2 の) エルミート保型形式へのリフトである．これを対角行列のなすエルミート上半空間の部分多様体に制限することを考える．これによって得られる周期積分を用いて，次数 4 のある L 関数の中心値の公式が得られたので，それについてお話ししたい．また，この制限の問題は Gross-Prasad 予想の状況ではないことを踏まえつつ， Gross-Prasad 予想の具体例となる結果との比較についても話したい．

野口和範 small category の分岐被覆とゼータ関数

この講演では small category の分岐被覆を紹介し，その性質について述べる．特に有限分岐被覆 P : E → B に対し， B のゼータが E のゼータを割り切ることを示す．これはデデキンド予想と呼ばれる代数体の拡大とそれらのデデキンドゼータの関係を示す予想の圏論的類似である．最後に素数定理やリーマン予想があるトポロジー問題と同値であるというビョルナーの私語をを紹介し， small category のゼータとオイラー標数，分岐被覆とのつながりについて述べる．

Taekyung Kim Descent on elliptic curves

In this talk, I will introduce descent on elliptic curves. After reviewing some basic stuffs briefly, I will give various interpretations of n-Selmer groups of elliptic curves.

町出智也 On Tornheim double series and generators for vector spaces spanned by double zeta values of even weight

重さが偶数の 2 重ゼータ値が生成するベクトル空間の次元の上限は，モジュラー関数の空間の次元を使って記述できることが知られている．本講演では，その上限と等しい個数の生成元の組を， Tornheim 2 重級数を使って具体的に構成する．

2 月 16 日 (日)

山田智宏 Explicit Chen’s theorem

We show that every even number greater than exp exp 24 can be represented as the sum of a prime and the product of at most two primes.

閔正媛有限群の Witten ゼータ関数と Witten L-関数について

ウィッテンゼータ関数とは，コンパクト位相群 G について，

ζ_{G}^{W} (s) =^{def} ∑_{ρ∈Ĝ}(deg ρ)^{−s}

と定義されるものであり，またウィッテン L-関数とは， g ∈ G について，

ζ_{G}^{W} (s; g) =^{def} ∑_{ρ∈Ĝ} \frac{χ(g)}{deg ρ}(deg ρ)^{−s}, χ(g) = trace(ρ(g))

と定義されるものである．ただし，

Ĝ = {G の既約ユニタリ表現} / ∼

である．有限群のウィッテン L-関数は，次より， g ≠ e のとき， ζ_{G}^{W} (−2 ; g) = 0 となる性質をもつ；

ζ_{G}^{W} (−2 ; g) = ∑_{ρ∈Ĝ} χ(g) deg ρ

= ∑_{ρ∈Ĝ} trace (ρ(g)) trace (ρ(e))

= ∑_{ρ∈Ĝ} trace (ρ(g)) \overline{trace (ρ(e))}

= 0

今回の講演では，有限群 S_{n}, B(m, n), S_{n} × ⋯ × S_{n} の場合について， s =?2 での位数について考える． (ただし， B(m, n) = C_{m}^{n} ⋊ S_{n}, C_{m} は位数 m の巡回群． )

赤塚広隆臨界領域内におけるオイラー積の挙動と素数分布について

オイラー積の収束は，素数分布，ゼータ関数の零点分布と密接に関係しており，非常に重要である．リーマンゼータ関数の場合， s = 1 に極があるため，実部が 1 より小のところではオイラー積は発散してしまう．しかし，極の寄与を適切に見積もることで，実部が 1 より小のところでのオイラー積の漸近挙動と，素数分布，零点分布との関係を調べることができる．本講演では，リーマン予想を仮定すると，臨界領域内のオイラー積の挙動は，リーマンゼータ関数の臨界領域内の値と von Mangoldt 関数の和を用いてかなり精密に書けることを紹介する．